OpenStudy (anonymous):

in a lab, we had to spin a stopper around in a constant circle and calculate the period and acceleration, what could be a cause for error for this experiment (no human errors)

OpenStudy (anonymous):

what errors can you think of?

OpenStudy (anonymous):

if your openion is "pandolm" so:T=1/2pi sqrt(l/g)-----pi number should use completly (its not possible because pi=3.141592...)-l=lenght of hanged stopper and there is not any instrument to measure excalty fact!-and g=gravity=9.80665... ithink that it depend on formula and the other thing is weight of stopper and acceloration of it depend on geometry and position on the ground.g=9.8 m/s^2 is an average -for exapmle in some places are more and in some places are less---the other thing is antisymmetric of shape that effect on center mass and ...